OpenStudy (anonymous):

If X is a uniformly distribution discrete random variable with possible values the integers from -3 to 5(inclusive). A) What is the probability distribution function for x? B)What is the probability that X is between 1 and 3 (inclusive) c) find the expect value of X d) find the variance of X (just use general computational formula) e) find the standard Deviation of X.

OpenStudy (kropot72):

A) Let r = -3, -2, -1, 0, 1, 2, 3, 4, 5. \[\large P(X = r)=\frac{1}{5-(-3)+1}=\frac{1}{9}\]

OpenStudy (anonymous):

Two ways of doing part (b). Use the CDF (once you find it) or add the appropriate probabilities.

OpenStudy (anonymous):

k i guess i will use the cdf

OpenStudy (anonymous):

1<=x<=3

OpenStudy (anonymous):

It's up to you, but the second method is much less work. Finding the CDF ends up being the same process in the end (if you recall that recent question you asked about finding the CDF of a discrete uniform distribution).

OpenStudy (anonymous):

yeah i recall that

OpenStudy (anonymous):

i guess the prof doesnt mind about the second way

OpenStudy (kropot72):

c) The expected value of X is found from: \[\large E(X)=\frac{1}{9}(-3-2-1+1+2+3+4+5)=you\ can\ calculate\]

OpenStudy (anonymous):

k isee

OpenStudy (anonymous):

i guess part b , d and e are left

OpenStudy (kropot72):

d) \[\large Var(X)=\sum_{}^{}x^{2}.\frac{1}{9}-[E(X)]^{2}\ for\ x=-3,-2,.........4,5 \]

OpenStudy (anonymous):

i see var(x)=summation (-3)^2*1/9- {E(x)]^2 know rest

OpenStudy (anonymous):

Now the last part is e)

OpenStudy (kropot72):

e) Take the square root of the variance.

OpenStudy (anonymous):

k thx so much @kropot72

OpenStudy (kropot72):

You're welcome :)